OpenStudy (anonymous):

find dy/dx for y=((2x+1)^1/2))(x^3)

OpenStudy (anonymous):

dy = 1/2 {(2x+1)^-1/2} * 3 x^2 dx dy/dx = 3/2 {x^2/(2x+1)^-1/2}